A ball of mass 0.5kg is dropped from a height of 0.8m unto a concrete floor and rebounds to a height of 0.5m. calculate the impulse of the floor on the ball
1
Expert's answer
2021-07-14T10:16:32-0400

Given:

"m=0.5\;\rm kg"

"h_1=0.8\;\rm m"

"h_2=0.5\;\rm m"

The impulse


"\Delta p=m|v_f-v_i|=m|\sqrt{2gh_f}-\sqrt{2gh_i}|"

"=0.5|\sqrt{2*9.8*0.5}-\sqrt{2*9.8*0.8}|=0.41\: \rm kg*m/s"